OpenStudy (anonymous):

-5(k + 4) >= 3(k - 4) A. k < = -1 B. k >= 32 C. k < = 12/5 D. k < = -5 E. k >= 123 F. k < = 8

OpenStudy (saifoo.khan):

A for Apple.

OpenStudy (anonymous):

-5k-20>=3k-12 -5k-3k>=-12+20 -k(5+3)>=8 -k>=8/(5+3) -k>=1 k<=-1

OpenStudy (anonymous):